what is the x and the y-intercepts of 3x-y=6
Answer by fcabanski(889)

The x-intercept is the place where the line crosses the x-axis. In other words, y=0. To find it, substitute 0 for y, solve for x. Then you have the x and y coordinates of the x-intercept.


3x-0=6


divide both sides by 3


x=2


The x-intercept is (2,0)


For the y-intercept set x=0, then solve for y. You can also manipulate the equation into the slope-intercept form (y=mx+b) where b is the y-coordinate of the y-intercept. This equation is set up in a way that it's easier to substitute 0 for x.


3(0)-y=6


-y=6


Divide both sides by -1.


y = -6


The y-intercept is (0,-6).

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com